OpenStudy (anonymous):

if the magnetic field is 2.1 m from a current carrying wire is 7.1x10^-5 T, what is the current flowing through the wire?

OpenStudy (anonymous):

If we assume the wire to be infinitesimal, we have, by Ampère's law, that: \[ \int_\gamma B\cdot dl=|B|2\pi r=\mu_0I \]Then: \[ |B|=\frac{\mu_0 I}{2\pi r} \]Solve for \(I\), the current in the enclosed wire.

OpenStudy (anonymous):

thanks! Does mu have a certain value?

OpenStudy (anonymous):

Indeed. \(\mu_0\) is the magnetic permeability of free space, which is, by definition: \[ \mu_0 = 4\pi \cdot 10^{-7} H/m \]

OpenStudy (anonymous):

And, sure thing.

OpenStudy (anonymous):

thanks a lot!